# Zagreb, Croatia Ensembling uncertainty with negative objectness

### v3

@@ 1 Introduction
Discriminative models may become robust to outliers through upgrade to open-set recognition [Liang et al.(2022)Liang, Wang, Miao, and Yang]. This can be conveniently achieved by complementing standard classification with outlier detection [Zhang et al.(2020)Zhang, Li, Guo, and Guo]. Typically, the outlier detector delivers a scalar outlier score that induces ranking and enables detection through thresholding [Ruff et al.(2021)Ruff, Kauffmann, Vandermeulen, Montavon, Samek, Kloft, Dietterich, and Müller]. Many outlier detectors both in the image-wide [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] and the pixel-level context [Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić, Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro] rely on negative training data. Although it cannot represent the entire variety of the visual world, the negative data can still help by signaling that not all data should be confidently recognized. If the variety of the negative data exceeds the variety of the inliers, then there is a reasonable hope that the test outliers will be detected. Unfortunately, this entails undesired bias towards test outliers that appear similar to the negative training data. This concern can be addressed either by relying on synthetic negatives [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li], or through separate ranking with respect to approaches that do not train on real negative data [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al., Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann].
@@
Many existing methods improve open-set performance of existing closed-set models by encouraging low-confidence predictions in negative training data [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song, Lee et al.(2018)Lee, Lee, Lee, and Shin]. However, this approach conflates prediction uncertainty with outlier recognition. In this paper, we disentangle the two factors by reconsidering the K+1-way classifier where the additional class represents the training negative data. This allows us to formulate the outlier score as the posterior of the K+1-th class, which we term negative objectness. Moreover, the outlier score can also be formulated as prediction uncertainty over the K known classes [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. We find that these two formulations exhibit complementary behaviour due to being sensitive to different test outliers. Consequently, we propose a novel outlier score that we term UNO: ensemble of Uncertainty over K inlier classes [Hendrycks and Gimpel(2017)] and Negative Objectness.
Our UNO score is a strikingly good fit in the pixel-level context as an extension of the direct set-prediction approach [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum]. We express the negative objectness as the posterior of the negative class in the frame of a K+2-way mask-level classifier. A closer look suggests that our score performs well due to different inductive bias of the two components as indicated by a remarkably weak correlation. The UNO score performs competitively in the image-wide context as well, even though the absence of the no-object class decreases the synergy of the two components. Our models outperform the state-of-the-art on several standard benchmarks in both learning setups, i.e. with and without real negative data.
@@ 2 Related work
Image-wide classification. Image-wide classification models can be complemented with outlier detection either through a modified training procedure or in a post-hoc manner with frozen parameters [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. Post-hoc methods build upon pre-trained closed-set classifiers. Early baselines express the outlier score as maximum softmax probability [Hendrycks and Gimpel(2016)] or maximum logit [Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] calibrates softmax probabilities with temperature scaling while ODIN [Liang et al.(2017)Liang, Li, and Srikant] pre-processes the input with anti-adversarial perturbations. Recent post-hoc methods simplify layer activations [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] or consider distances from class prototypes [Zhang et al.(2022)Zhang, Fu, Chen, Du, Li, Wang, Han, Zhang, et al.]. In contrast, training methods involve various regularizations that either consider only inlier data or utilize real negative data as well. The former rely on modeling confidence [DeVries and Taylor(2018)], self-supervised training [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song, Tack et al.(2020)Tack, Mo, Jeong, and Shin], per-sample temperature scaling [Hsu et al.(2020)Hsu, Shen, Jin, and Kira], outlier synthesis in feature space [Du et al.(2022)Du, Wang, Cai, and Li, Tao et al.(2023)Tao, Du, Zhu, and Li] or mitigating overconfidence [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li]. Recent methods guide the learning process with respect to nearest neighbours [Park et al.(2023)Park, Jung, and Teoh] or leverage distances in feature space either by combining different distance functions [Noh et al.(2023)Noh, Jeong, and Lee] or distances in different feature spaces [Chen et al.(2023)Chen, Lin, Xu, and Vela]. Early work in training with negative data [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] encourages classifiers to produce high entropy in inlier samples. MCD [Yu and Aizawa(2019)] promotes discrepancy between prediction on OOD samples of two classification heads. UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] groups negative data following inlier taxonomy to enrich semantic knowledge of inlier classes. MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] interpolates between inliers and negatives to improve regularization. Unlike related training based methods UNO does not modify the standard classifier training objective, but merely requires a slight architectural change to extend the classifier with a logit for the negative class. Additionally, UNO requires significantly less negative data than related methods.
@@
Dense prediction. Many image-wide outlier detection methods can directly complement models for pixel-level prediction [Hendrycks and Gimpel(2016), Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. Early approaches estimate the prediction uncertainty with maximum softmax probability [Hendrycks and Gimpel(2017)], ensembling [Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell] or Bayesian uncertainty [Mukhoti and Gal(2018)]. Subsequent work suggests using energy scores [Liu et al.(2020)Liu, Wang, Owens, and Li] or standardized max-logit [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo]. Training with negative data can be incorporated by pasting negative content from broad negative datasets (usually ADE20K [Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba] or COCO [Lin et al.(2014a)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]) atop inlier images [Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro, Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić]. However, real negative data can be replaced with synthetic negative patches obtained by sampling a jointly trained generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Zhao et al.(2021)Zhao, Cao, and Lin]. Outlier detection performance can be further improved by providing more capacity [Vaze et al.(2021)Vaze, Han, Vedaldi, and Zisserman]. Another branch of outlier detection methods considers generative models [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Liang et al.(2022)Liang, Wang, Miao, and Yang].
@@
Mask-level recognition Recent panoptic architectures [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum] decompose scene understanding into class-agnostic segmentation and region-wide recognition. They frame the detection of semantic masks as direct set-prediction where each mask is classified into K inlier classes and one no-object class. This induces a degree of openness and therefore favours outlier detection. Recent work shows great applicability of the Mask2Former architecture for open-set segmentation [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu, Grcić et al.(2023)Grcić, Šarić, and Šegvić]. All these methods detect anomalies due to not belonging to any inlier class and improve when training with negative data. Specifically, RbA [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney] reduces energy in negative pixels, Mask2Anomaly [Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo] uses a contrastive loss, while EAM [Grcić et al.(2023)Grcić, Šarić, and Šegvić] ensembles region-wide outlier scores. However, these approaches conflate negative data with no-object regions. Furthermore, they fail to exploit the fact that mask queries behave like one-vs-all classifiers [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu], which makes them appropriate for direct detection of anomalous regions. UNO addresses these weaknesses by decoupling negatives from the no-object class, and representing them as a standalone class in the K+2-way taxonomy.
@@ 3 Disentangling negative objectness from uncertainty
Let $h_{\theta_{1}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a feature extractor (e.g. ResNet [He et al.(2016)He, Zhang, Ren, and Sun] or ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]) onto which we attach our K+1-way classifier $g_{\theta_{2}}:\mathbb{R}^{d}\rightarrow\Delta^{K}$, where $\Delta^{K}$ is a K-dimensional probabilistic simplex. Given an input $\mathbf{x}\in\mathcal{X}$, we compute a latent representation $\mathbf{z}=h_{\theta_{1}}(\mathbf{x})$ and map it onto the simplex point $\mathbf{p}=\text{softmax}(g_{\theta_{2}}(\mathbf{z}))$. Thus, our deep model $f_{\theta}=g_{\theta_{2}}\circ h_{\theta_{1}}$ maps input samples from $\mathcal{X}$ onto simplex points in $\Delta^{K}$. We assume that the classifier $g_{\theta_{2}}$ is a simple projection. Thus we have $g_{\theta_{2}}(\textbf{x})=\textbf{W}\cdot\textbf{x}$, where the rows of W correspond to class vectors $\textbf{w}_{j}$.
Negative objectness. Given feature representation $\mathbf{z}$ and a linear classifier $g$, we formulate our negative objectness score ($\textbf{s}_{\text{NO}}$) as posterior of the K+1-th class:
@@
Prediction uncertainty. Negative objectness is not suitable for detecting test outliers that are semantically dissimilar from the training data. However, many of these outliers yield uncertain predictions across inlier classes of our K+1-classifier. We address this observation by defining the uncertainty score as negative prediction confidence across the K inlier classes:
@@
UNO score. We integrate the prediction uncertainty (2) and negative objectness (1) into the UNO score as the sum of the two terms:
@@
UNO has an interesting geometrical interpretation in the pre-logit space $\mathbb{R}^{d}$. We note that class vectors tend to be mutually orthogonal: $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$. This arrangement is enforced by the standard supervised loss ($-\log\text{softmax}$) whenever there is enough data, as well as enough capacity and dimensionality ($d>K+1$) in the feature extractor $h_{\theta_{1}}$. Thus, a small angle between the latent representation z and the negative vector $\textbf{w}_{K+1}$ leads to high outlier objectness $\textbf{s}_{\text{NO}}$. Moreover, we note that the uncertainty score $\textbf{s}_{\text{Unc}}$ is negatively correlated with the feature norm $\lvert\lvert\mathbf{z}\rvert\rvert$. If we fix the direction of z, its norm can be viewed as reciprocal softmax temperature. As usual, large temperatures lead to uncertain predictions [Hinton et al.(2015)Hinton, Vinyals, and Dean].
The above insights show that the two components of our score will be at least partially decorrelated due to being sensitive to different outliers. Thus, the UNO score is likely to achieve lower error than any of the two components alone. Figure 2 illustrates our insight in an image-wide recognition experiment on small images. Further empirical analysis indicates a weak positive correlation between the UNO components, as detailed in the Appendix.
@@
Extending UNO for dense prediction. The UNO outlier detector can be attached to any pre-trained deep classifier. In the case of image-wide prediction, we simply extend a pre-trained K-way classifier with an additional class. However, in the case of dense prediction, many contemporary architectures rely on mask-level recognition [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar]. Consequently, we build our per-pixel score by merging multiple mask-level UNO scores:
@@
Equation 4 assumes that a dense prediction model has partitioned the input image into $N$ different masks $\mathbf{m}_{i}$ with the corresponding latent mask embeddings $\mathbf{z}_{i}$.

### v4

@@ 1 Introduction
Discriminative models may become robust to outliers through upgrade to open-set recognition [Liang et al.(2022)Liang, Wang, Miao, and Yang]. This can be conveniently achieved by complementing standard classification with outlier detection [Zhang et al.(2020)Zhang, Li, Guo, and Guo]. Typically, the outlier detector delivers a scalar outlier score that induces ranking and enables detection through thresholding [Ruff et al.(2021)Ruff, Kauffmann, Vandermeulen, Montavon, Samek, Kloft, Dietterich, and Müller]. Many outlier detectors both in the image-wide [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] and the pixel-level context [Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić, Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro] rely on negative training data. Although it cannot represent the entire variety of the visual world, the negative data can still help by signaling that not all data should be confidently recognized. If the variety of the negative data greatly exceeds the variety of the inliers, then there is a reasonable hope that the test outliers will be detected. Unfortunately, this entails undesired bias towards test outliers that appear similar to the negative training data. This concern can be addressed either by relying on synthetic negatives [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Grcić and Šegvić(2024b)], or through separate ranking with respect to approaches that do not train on real negative data [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al., Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann].
@@
Many existing methods improve open-set performance of existing closed-set models by encouraging low-confidence predictions in negative training data [Hendrycks et al.(2022a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song, Lee et al.(2018a)Lee, Lee, Lee, and Shin]. However, this approach conflates prediction uncertainty with outlier recognition. In this paper, we disentangle the two factors by reconsidering the K+1-way classifier where the additional class represents the negative data. This allows us to formulate the outlier score as the posterior of the K+1-th class, which we term negative objectness. Moreover, the outlier score can also be formulated as prediction uncertainty over the K known classes [Hendrycks et al.(2022a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. We find that these two formulations exhibit complementary behaviour due to being sensitive to different test outliers. Consequently, we propose a novel outlier score that we term UNO: ensemble of Uncertainty over K inlier classes [Hendrycks and Gimpel(2017)] and Negative Objectness.
Our UNO score is a strikingly good fit in the pixel-level context as an extension of some direct set-prediction approach [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko]. We express the negative objectness as the posterior of the negative class in the frame of a K+2-way mask-level classifier [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum]. A closer look suggests that our score performs well due to different inductive bias of the two components as indicated by a remarkably weak correlation. The UNO score performs competitively in the image-wide context as well, even though the absence of the no-object class decreases the synergy of the two components. Our models outperform the state-of-the-art on several standard benchmarks in both learning setups, i.e. with and without real negative data.
@@ 2 Related work
Image-wide classification models can be complemented with outlier detection either through a modified training procedure or in a post-hoc manner with frozen parameters [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. Post-hoc methods build upon pre-trained closed-set classifiers. Early baselines express the outlier score as maximum softmax probability [Hendrycks and Gimpel(2016)] or maximum logit [Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] calibrates softmax probabilities with temperature scaling while ODIN [Liang et al.(2017)Liang, Li, and Srikant] pre-processes the input with anti-adversarial perturbations. Recent post-hoc methods simplify layer activations [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] or fit generative models to pre-logit features [Zhang et al.(2022)Zhang, Fu, Chen, Du, Li, Wang, Han, Zhang, et al., Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Taghikhah et al.(2024)Taghikhah, Kumar, Segvic, Eslami, and Gumhold]. In contrast, training methods involve various regularizations that either consider only inlier data or utilize real negative data as well. The former rely on modeling confidence [DeVries and Taylor(2018)], self-supervised training [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song, Tack et al.(2020)Tack, Mo, Jeong, and Shin], per-sample temperature scaling [Hsu et al.(2020)Hsu, Shen, Jin, and Kira], outlier synthesis in feature space [Du et al.(2022)Du, Wang, Cai, and Li, Tao et al.(2023)Tao, Du, Zhu, and Li, Kumar et al.(2023)Kumar, Segvic, Eslami, and Gumhold] or mitigating overconfidence [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li]. Recent methods guide the learning process with respect to nearest neighbours [Park et al.(2023)Park, Jung, and Teoh] or leverage distances in feature space either by combining different distance functions [Noh et al.(2023)Noh, Jeong, and Lee] or distances in different feature spaces [Chen et al.(2023)Chen, Lin, Xu, and Vela]. Early work in training with negative data [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] encourages high entropy in inlier samples. MCD [Yu and Aizawa(2019)] promotes discrepancy between prediction on OOD samples of two classification heads. UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] groups negative data following inlier taxonomy to enrich semantic knowledge of inlier classes. MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] interpolates between inliers and negatives to improve regularization. Unlike related training-based methods, UNO does not modify the standard classifier training objective, but merely requires a slight architectural change to extend the classifier with a logit for the negative class. Moreover, empirical insights suggest that UNO requires significantly less negative data than related methods.
Dense prediction.
@@
Many image-wide outlier detection methods can directly complement models for pixel-level prediction [Hendrycks and Gimpel(2016), Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. Early approaches estimate the prediction uncertainty with maximum softmax probability [Hendrycks and Gimpel(2017)], ensembling [Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell] or Bayesian uncertainty [Mukhoti and Gal(2018)]. Subsequent work suggests energy scores [Liu et al.(2020)Liu, Wang, Owens, and Li] and the standardized max-logit score [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo]. Training with negative data can be incorporated by pasting negative content from broad negative datasets (ImageNet [Bevandić et al.(2019)Bevandić, Krešo, Oršić, and Šegvić], ADE20K [Zhou et al.(2017b)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba] or COCO [Lin et al.(2014a)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]) atop inlier images [Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro, Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić]. However, real negative data can be replaced with synthetic negative patches obtained by sampling a jointly trained generative model [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Zhao et al.(2021)Zhao, Cao, and Lin]. Outlier detection performance can be further improved by providing more capacity [Vaze et al.(2021)Vaze, Han, Vedaldi, and Zisserman]. Another branch of methods considers generative models [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Liang et al.(2022)Liang, Wang, Miao, and Yang].
Mask-level recognition.
@@
Recent panoptic architectures [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum] decompose scene understanding into class-agnostic segmentation and region-wide recognition. They frame the detection of semantic masks as direct set-prediction where each mask is classified into K inlier classes and one no-object class. This induces a degree of openness and therefore favours outlier detection. Recent work shows great applicability of the Mask2Former architecture for open-set segmentation [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu, Grcić et al.(2023)Grcić, Šarić, and Šegvić]. All these methods detect anomalies due to not belonging to any inlier class and improve when training with negative data. Specifically, RbA [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney] reduces energy in negative pixels, Mask2Anomaly [Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo] uses a contrastive loss, while EAM [Grcić et al.(2023)Grcić, Šarić, and Šegvić] ensembles region-wide outlier scores. However, these approaches conflate negative data with no-object regions. Furthermore, they fail to exploit the fact that mask queries behave like one-vs-all classifiers [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu], which makes them appropriate for direct detection of anomalous regions. UNO addresses these weaknesses by decoupling negatives from the no-object class, and representing them as a standalone class in the K+2-way taxonomy.
@@ 3 Disentangling negative objectness from uncertainty
Let $h_{\theta_{1}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a feature extractor (e.g. ResNet [He et al.(2016)He, Zhang, Ren, and Sun] or ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]) onto which we attach our K+1-way classifier $g_{\theta_{2}}:\mathbb{R}^{d}\rightarrow\Delta^{K}$, where $\mathcal{X}$ is the input space and $\Delta^{K}$ is a K-dimensional probabilistic simplex. Given an input $\mathbf{x}\in\mathcal{X}$, we compute a latent representation $\mathbf{z}=h_{\theta_{1}}(\mathbf{x})$ and map it onto the simplex point $\mathbf{p}=\text{softmax}(g_{\theta_{2}}(\mathbf{z}))$. Thus, our deep model $f_{\theta}=g_{\theta_{2}}\circ h_{\theta_{1}}$ maps input samples from $\mathcal{X}$ onto simplex points in $\Delta^{K}$. We assume that the classifier $g_{\theta_{2}}$ is a simple projection. Thus we have $g_{\theta_{2}}(\textbf{x})=\textbf{W}\cdot\textbf{x}$, where the rows of W correspond to class vectors $\textbf{w}_{j}$.
Negative objectness.
Given feature representation $\mathbf{z}$ and a linear classifier $g$, we formulate the negative objectness score ($\textbf{s}_{\text{NO}}$) as posterior of the K+1-th class:
@@
Prediction uncertainty.
Negative objectness is not suitable for detecting test outliers that are semantically dissimilar from the negative training data. However, many of these outliers yield uncertain predictions across inlier classes of our K+1-classifier. We address this observation by defining the uncertainty score as negative prediction confidence across the K inlier classes:
@@
UNO score.
We integrate the prediction uncertainty (2) and negative objectness (1) into the UNO score as the sum of the two terms:
@@
UNO has an interesting geometrical interpretation in the pre-logit space $\mathbb{R}^{d}$. Class vectors tend to be mutually orthogonal: $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$ [Wang et al.(2022)Wang, Li, Feng, and Zhang]. This arrangement is enforced by the standard supervised loss ($-\log\text{softmax}$) whenever there is enough data, and enough capacity and dimensionality ($d>K+1$) in the feature extractor $h_{\theta_{1}}$. Thus, a small angle between the latent representation z and the negative vector $\textbf{w}_{K+1}$ leads to high outlier objectness $\textbf{s}_{\text{NO}}$. On the other hand, the uncertainty score $\textbf{s}_{\text{Unc}}$ is negatively correlated with the feature norm $\lvert\lvert\mathbf{z}\rvert\rvert$. If we fix the direction of z, its norm can be viewed as reciprocal softmax temperature. As usual, large temperatures lead to uncertain predictions [Hinton et al.(2015)Hinton, Vinyals, and Dean].
The above insights show that the two components of our score will be at least partially decorrelated due to being sensitive to different outliers: $\textbf{s}_{\text{NO}}$ detects large feature norms and good alignment with $\mathbf{w}_{K+1}$, while $\textbf{s}_{\text{Unc}}$ detects small feature norms and poor alignment with inlier vectors. Thus, the UNO score is likely to achieve lower error than any of the two components alone. Figure 2 illustrates our insight in an image-wide recognition experiment on small images. Further empirical analysis indicates a weak positive correlation between the UNO components, as detailed in the Appendix.
@@
Extending UNO for dense prediction.
The UNO outlier detector can be attached to any pre-trained deep classifier. In the case of image-wide prediction, we simply extend a pre-trained K-way classifier with an additional class. However, in the case of dense prediction, many contemporary architectures rely on mask-level recognition [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar]. Consequently, we build our per-pixel score by averaging the mask-level UNO scores:
@@
Note that equation 4 assumes that the dense prediction model has partitioned the input image into $N$ different masks $\mathbf{m}_{i}$ with the corresponding latent mask embeddings $\mathbf{z}_{i}$.